OpenStudy (anonymous):

1 + 1 = 3 o . o 2 + 2 = 7 3 + 3 = ?

OpenStudy (anonymous):

3 +3 = 13

OpenStudy (anonymous):

I think 11

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

see 1+1=3 but 1+1= 2 diff in result =1 similarly diif in 2 is 3 so i suggest diff in 3rd should be 5 3+3+5=11

OpenStudy (anonymous):

You beat me in that man

OpenStudy (anonymous):

:)

OpenStudy (saifoo.khan):

11

OpenStudy (anonymous):

right

OpenStudy (saifoo.khan):

Like always!